y=3x+1 i need to plot this on a graph and then make a straight line my graph is y 10 to 0 then to -3 and X is -5 to 0 then 0 to 6
Answer by Edwin McCravy(20055)


The only points with integer coordinates that will 
fit on your graph are these 5

 x | y = 3x+1
----------------------------
-1 | y = 3(-1)+1 = -3+1 = -2
 0 | y = 3(0)+1 = 0+1 = 1
 1 | y = 3(1)+1 = 3+1 = 4
 2 | y = 3(2)+1 = 6+1 = 7
 3 | y = 3(3)+1 = 9+1 = 10

So we plot (-1,-2), (0,1), (1,4), (2,7), (3,10)
and draw a line through them.



Edwin